OpenStudy (anonymous):

Solve. (X-9)^2 = x^2

OpenStudy (anonymous):

x^2-18x-81=x^2 -18x=-81 x=4.5

OpenStudy (anonymous):

im having problems with this question because i dont know what it wants me to do Research the highest interest rate (APY—annual percent yield) for 2-year and 5-year CDs. Document the company's name, interest rate, and minimum investment. The minimum investment must be less than or equal to $5,000.

OpenStudy (kirbykirby):

@pandaluver44 \((x-9)^2 \neq x^2-9^2 \)

OpenStudy (anonymous):

really? it doesn't?

OpenStudy (kirbykirby):

No, do not get it confused with the identity: \((x-9)(x+9)=x^2-9^2\)

OpenStudy (anonymous):

Would x= 9/2 be correct?

OpenStudy (anonymous):

oh. oops. sorry about that. I'll delete my posts.

OpenStudy (kirbykirby):

@Nancyo yes

OpenStudy (anonymous):

Yeah. Thanks